Hi, please answer this with an explanation. Thank

JP, Inc. has adopted the double-declining balance depreciation method for all of its equipment with half-year convention.

The company purchased equipment for $15,000 on August 1, 2011. This equipment has an estimated life of 8 years and an expected salvage value of $3,000.

This asset is sold at the end of February 2013 for $9,000 in cash. How much gain or loss does JP, Inc. report on its 2013 Income Statement?
Answer #1

Double declining Method

A

Cost

$       15,000.00

B

Residual Value

$         3,000.00

C=A - B

Depreciable base

$       12,000.00

D

Life [in years]

8

E=C/D

Annual SLM depreciation

$         1,500.00

F=E/C

SLM Rate

12.50%

G=F x 2

DDB Rate

25.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

2011

$    15,000.00

12.50%*

$    1,875.00

$ 13,125.00

$            1,875.00

2012

$    13,125.00

25.00%

$    3,281.25

$    9,843.75

$            5,156.25

2013

$       9,843.75

12.50%

$    1,230.47

$    8,613.28

$            6,386.72

*Double declining balance uses double of SLM rate but since half year convention is applied only hald of the depreciation will be charged in 2011 and 2013.

Calculation of profit and loss on disposal.

Cost

$ 15,000.00

Accumulated depreciation

$    6,386.72

Book value

$    8,613.28

Sales price

$    9,000.00

Book value

$    8,613.28

Gain /(loss)

$        386.72

Profit on disposal = $386.72 or $389